A binomial probability experiment is conducted with the given parameters. Compute the probability of x successes in the n independent trials of the experiment. n equals 9​, p equals 0.9​, x less than or equals 3
Answer #1

Answer

We have sample size n= 9

probability p = 0.9

we have to find probability of less than or equal 3 trials.

Using the binomial probability formula, we can get the probability of getting 0,1,2 and 3

P(x = 0) = *pr * (1-p)-T 9 0 ) * 0.90 * (1-0.9)9-0ー0.00000000

P(x = 1) = inom{n}{r}*p^r*(1-p)^{n-r} = inom{9}{1}*0.9^1*(1-0.9)^{9-1} = 0.00000000

P(x = 2) = 9 *pr * (1-p)-T 2 ) * 0.92 * (1-0.9)9-2 = 0.00000292

P(x = 3) = 9 *pr * (1-p)-r * 0.9* * (1 - 0.9)9-3

Adding all the probabilities, we get

Probability of less than or equal to 3 = P(x=0)+P(x+1)+P(x=2)+P(x=3) = 0.00000000+0.00000000+0.00000292+0.00006124

we get

P(x less than or equal to 3) = 0.00006423